If a jar contains 23 red marbles and 4 green marbles, what is the probability that a green marble will be selected at random?
Answer by checkley77(12569)

Probability=success/total outcomes.
P=4/(23+4)=4/27=.148148 or 14.8% chance of drawing a green marble